16. Mrs. Bailey gives a test, and her students’ scores range from 30 to 70. She decides to curve the scores, so that they range from 65 to 95. Let “x” be an original score, and “y” be a curved score. Using the ordered pairs (30,65) and (70,95)
a. In the scenario given above, what will an original score of 62 become?

Found 3 solutions by ikleyn, ewatrrr, greenestamps:
Answer by ikleyn(52798)

.

. . . to curve the scores . . .


???????????????????????????????????????????


        Please do not scare me . . .



Answer by ewatrrr(24785)

 

 Using the ordered pairs:
30,65) and 
(70,95)  m = -30/-40 = 3/4

Let y represent the converted score x
 y - 65 = (3/4)x -(3/4)30
 y = (3/4)x  + 260/4 - (3/4)30 = (3/4)x + 170/4

a. In the scenario given above, what will an original score of 62 become?

 y = (3/4)x + 170/4  = (3/4)62 + 170/4 = 186/4 + 170/4 = 356/4 = 89



Answer by greenestamps(13200)



An informal solution, without algebra....

With the actual scores ranging from 30 to 70, the score of 62 is (62-30)/(70-30) = 32/40 = 4/5 of the way from 30 to 70.

In the new range of scores from 65 to 95, the new score corresponding to an original score of 62 should be 4/5 of the way from 65 to 95.

95-65 = 30
(4/5)(30) = 24
65+24 = 89

ANSWER: 89